OpenStudy (sonyalee77):

The quotient of a number and four is twelve less than the number. What is the number? A. –4 B. 9.6 C. 8 D. 16

OpenStudy (austinl):

\(\displaystyle \frac{x}{4}=x-12\) Solve for x.

OpenStudy (sonyalee77):

k thanx :)

OpenStudy (austinl):

No problemo!

OpenStudy (sonyalee77):

is it 16?

OpenStudy (austinl):

It is indeed. :)

OpenStudy (sonyalee77):

YESSS! Thanx so much ! :D

OpenStudy (austinl):

You're very welcome, good job!